OpenStudy (anonymous):

How do I find the range of y = sin(3x)^4 + 2

OpenStudy (anonymous):

Is this \(\sin((3x)^4)\) or \((\sin(3x))^4\)?

OpenStudy (anonymous):

(sin(3x))^4

OpenStudy (anonymous):

Either way, you can find the range by using the fact that \(-1\le \sin(z)\le 1\).

OpenStudy (anonymous):

did you try graphing it ?

OpenStudy (anonymous):

I'm not going to take the easy way out by graphing

OpenStudy (anonymous):

So the range of sin(x) i know is [-1,1], then you take those things to the 4th power to get 2 1's. What now?

OpenStudy (anonymous):

Okay. First we have \(-1\le \sin(3x) \le 1\). Now take the fourth power of both sides: \(0\le (\sin(3x))^4 \le 1 \implies 2\le (\sin(3x))^4+2\le 3.\)

OpenStudy (anonymous):

So the range of \(y\) is \([2,3]\).

OpenStudy (anonymous):

where did you get the 0?

OpenStudy (anonymous):

When you take an even power of an inequality you have to be careful with numbers of different sign. Consider for example \(-5<1 \) does not imply \(25<1\). So you first have to divide the inequality into two parts, from -1 to 0 and then from 0 to 1. Do it this way and you'll get what I did.

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

did you study derivatives ?

OpenStudy (anonymous):

No, I'm in algebra 2

OpenStudy (anonymous):

oh...ok